if a triangle has a base of (4x-2) inches and a height of 3x inches, what is the area in the simplified form?

Answer by Cromlix(4381)

area of a triangle = 1/2(base * height)
area of a triangle = 1/2 ((4x - 2)(3x))
area of a triangle = 1/2 (12x^2 - 6x)
Area in simplified form = 6x^2 - 3x
Hope this helps.
:-)